From the edge of the top of the building with a height of 30.0m, Erica kicks a soccer ball with a velocity of 12 m/s horizontally.
A. When will the soccer ball strike the ground?
B. What is the velocity of the soccer ball when it strikes the ground?
1
Expert's answer
2021-10-15T10:23:41-0400

"A. \ t=\sqrt{2h/g}=\sqrt{2\cdot 30/9.8}=2.47\ (s)"


"B.\ v=\sqrt{2gh}=\sqrt{2\cdot 9.8\cdot 30}=24.2\ (m/s)"